# Quantum-thermodynamic enhancements in continuous thermal machines require energetic coherence

### v1

@@ Abstract
Quantum coherence has been shown to impact the operational capabilities of quantum systems performing thermodynamic tasks in a significant way, and yet the possibility of genuine coherence-enhanced thermodynamic operation remains unclear. Here we show that only the presence of energetic coherence —coherence between levels with different energies— in steady-state quantum thermal machines can lead to genuine thermodynamic advantage. On the other hand, engines showing coherence between degenerate levels, or subjected to noise-induced coherence, are shown to be systematically outperformed by classical stochastic engines using exactly the same set of (incoherent) resources. We illustrate our results with three prototypical models of heat engines and refrigerators and employ multi-objective optimization techniques to characterize quantum-enhanced regimes in connection with the thermodynamic uncertainty relation.
@@ I Introduction
Figure 1: Schematic representation of (a) the coherent three-level amplifier coupled to baths at hot $\beta_{\mathrm{h}}$ and cold $\beta_{\mathrm{c}}$ temperatures and coherent external driving, (b) the three-qubit autonomous (absorption) refrigerator where each qubit is locally coupled to baths at hot, cold, and intermediate temperatures $\beta_{\mathrm{m}}$, and (c) the noise-induced-coherence machine showing collective jumps induced by the baths at hot and cold temperatures, together with a classical work source given by an infinite-temperature bath.
@@
In this paper we show how to identify genuine coherence-induced quantum thermodynamic advantage and how to quantitatively characterize it in steady-state quantum thermal machines by combining three main ingredients. The first one is the systematic construction of thermodynamically-equivalent classical thermal machines that are able to produce the same average currents than their quantum-coherent counterparts, while using exactly the same amount of incoherent resources (essentially bath temperatures and energy structure). However, even if same average currents can be reproduced by classical models, fluctuations as captured by the variance of the currents might present significant differences. The comparison of fluctuations in the output currents (reliability of the machine) with respect to the classical counterpart is henceforth the second necessary ingredient in our analysis. This is connected with violations in the so-called thermodynamic uncertainty relation (TUR) [24, 25, 26], which provides a universal trade-off relation between power, efficiency, and output reliability for any classical Markovian thermal machine operating at steady-state conditions [27]. The TUR sets up a model-independent limit in the maximum reliability achievable by any classical machine at a given power output and efficiency. Hence the observation of TUR violations in quantum Markovian machines working in steady-state conditions [28, 29, 30, 31, 32, 33] may be considered as an unambigüous witness of a quantum thermodynamic signature. Nevertheless, the violation of the TUR in specific parameter regimes is not yet a sufficient (nor necessary) condition for ensuring a practical quantum-thermodynamic advantage in thermal machines, since these parameter regimes may be far from optimal performance. The third ingredient that we employ here is the use of multi-objective optimization techniques (Pareto optimization) [34, 35, 36, 37, 38, 39] to actually confirm enhanced stability (precision), power output or efficiency, with respect to classical thermodynamically-equivalent machines, in relevant (optimal) regimes of operation of the machine, such as maximum power conditions.
@@
Although our analysis is applicable to a broad class of steady-state quantum thermal machines, we employ three well-known (prototypical) models showing coherent-induced evolution (see Fig. 1), each corresponding to one of the three possible types of coherence that can arise in the working medium when approaching the steady state: (a) coherence between different energy levels (energetic coherence) induced by external driving, (b) coherence between levels with degenerate energies induced by internal Hamiltonian couplings, and (c) noise-induced coherence on degenerate levels induced by the reservoir. Our results show that only in the case of energetic coherence (case a) a practical quantum-thermodynamic advantage can be unambiguously identified, leading to enhanced optimal regimes not allowed by any equivalent classical engine using the same amount of thermodynamic resources. On the contrary, coherence in degenerate energy levels (cases b and c) can only lead to disadvantages in the performance, even in cases where the quantum machine dynamics contains intrinsic quantum features.
@@
This paper is organized as follows: Sec. II introduces the different thermal machine models in detail. Sec. III discusses how to characterize the thermodynamic performance in such systems. In Sec. IV, we describe the concept of classical equivalent machines and provide a general recipe for constructing them in the presence of different types of coherence. Sec. V.1 demonstrates that, in the presence of energetic coherence, quantum enhancements can be achieved with respect to classical equivalent machines. Sec.  V.2 demonstrate instead the disadvantages of quantum machines with non-energetic (i.e. degenerate) coherence compared to their classical counterparts. In Sec. VI, we optimize both quantum and classical thermal machines and compare their optimal solutions. Finally, Sec. VII provides a summary and conclusions of our main results.
@@ II Quantum thermal machines models
where $H(t)$ is given in Eq. (1) and $\mathcal{D}^{(r)}_{k}(\rho)$ denote the so-called dissipators, taking into account the effects of dissipative process $k$ from thermal reservoir $r$ on the system. These dissipators are given in terms of Lindblad operators $L^{(r)}_{k}$ associated to each reservoir:
@@
with $\alpha_{ij}^{k}=1$ if $\epsilon_{j}-\epsilon_{i}=\Delta\epsilon_{k}$ and 00 otherwise. In the case of non-degenerate transitions, the above operators reduce to simple jumps $L_{k}^{(r)}=\sqrt{\gamma_{ij}}\ket{j}\bra{i}$ | between energy levels $i\rightarrow j$. However, the operators in Eq. (4) can also describe collective jumps where two or more transitions with same (degenerate) gap $\Delta\epsilon_{k}$ in $H_{0}$ may occur simultaneously, e.g. $i\rightarrow j$ and $i\rightarrow j^{\prime}$ if $\epsilon_{j}=\epsilon_{j}^{\prime}$. In any case, every transition is connected to a single thermal bath, the rates $\gamma_{ij}\geq 0$ being time-independent and verifying the local detailed balance relation $\gamma_{ij}=\gamma_{ji}~{}e^{-\beta_{r}(\epsilon_{j}-\epsilon_{i})}$, with $\beta_{r}=1/{k_{B}T_{r}}$ the inverse temperature of the bath $r$ and $k_{B}$ Boltzmann’s constant. In the long-time limit, the evolution dictated by Eq. (3) converges to a steady-state, verifying $\mathcal{L}(\pi)=0$, where we denote $\pi(t)$ the steady state density operator. We notice that, due to the presence of the time-dependent Hamiltonian $V(t)$, the steady state can show a (residual) periodic time-dependence in the phase in the Schrödinger picture, which disappears when moving to a rotating frame.
Although the results we present here can be applied to any model of quantum thermal machine whose evolution can be described within the general framework introduced above, in this paper we will mainly focus on three representative and well-known models of quantum thermal machines, as illustrated in Fig. 1. They lead to three different types of coherent evolution respectively: the coherent three-level maser (Fig. 1a) which induces energetic coherence in the $H_{0}$ basis; the three-qubit autonomous quantum refrigerator (Fig. 1b), leading to Hamiltonian-induced coherence in a degenerate subspace of $H_{0}$, and the noise-induced-coherence engine (Fig. 1c), showing coherence between degenerate levels induced by collective bath transitions.
@@ II.3 Noise-induced-coherence machine
A notable example of this coherence is found in a 4-level absorption refrigerator, whose thermodynamics have been examined in previous studies [66, 69, 68]. In this system, two of the levels possess the same energy, as depicted in Fig. 1c. These levels are subjected to the influence of two distinct thermal baths, the action of both of which results in the emergence of horizontal coherences that persist in the system’s steady state. The system’s Hamiltonian in this case is as follows
@@ III Thermodynamic performance
where we recall that $\pi(t)$ may acquire a periodic time-dependence (in Schrödinger picture) due to the presence of non-diagonal elements (coherences) in the steady-state density operator. We also emphasize that for weak perturbations as the ones considered here, only the bare Hamiltonian $H_{0}$ enters in the heat currents [71, 72], ensuring consistency with the second law [73]. The first law takes the form $\langle\dot{W}\rangle=\sum_{r}\langle\dot{Q}_{r}\rangle$, imposing that any output power is sustained by input heat currents from the baths. Explicit expressions of the heat currents $\langle\dot{Q}_{r}\rangle$ valid for generic machines (with or without degeneracy) are given in Appendix A.
@@
where $\langle\dot{N}_{r}\rangle$ is the probability current associated to the reservoir $r$ (flux of quanta). It results from the fact that in the models baths are associated to a single energy gap $\Delta\epsilon_{r}$. In the steady state, due to the absence of leaks, the system will exchange excitations with all baths at the same rate, i.e. $\forall i,j$ we have $\langle\dot{N}_{i}\rangle=\langle\dot{N}_{j}\rangle:=\langle\dot{N}\rangle$. In addition, the first law constrains the average value of the work to satisfy $\langle\dot{W}\rangle=\sum_{r}\Delta\epsilon_{r}\langle\dot{N}\rangle=\omega_{d}\langle\dot{N}\rangle$. The ratio between the different steady-state currents hence verify
@@
By combining the above proportionality relations with the efficiency bounds above we can construct a table for the operational modes of each model (see Table 1).
@@
In any of the three cases, the TUR implies that beyond a certain threshold, a classical Markovian engine can only enhance its precision in the output (cooling) power at the cost of either reducing the output itself or reducing the energy efficiency, so that the above ratio remains bounded by $2$, i.e. $\mathcal{Q}\geq 2$. However, some models of quantum thermal machines have been shown to produce violations of the TUR, that is, they verify $\mathcal{Q}<2$ (see e.g. Refs. [28, 29, 30, 31, 32, 33]). Such violations act as a witness indicating an enhanced tradeoff between power, precision and efficiency that arise in certain parameter regimes. Nevertheless, it is in principle unclear whether that violations may occur in relevant regimes, i.e. where the machine performs in an optimal way, providing a practical quantum-thermodynamic advantage. In addition, such quantum-thermodynamic advantages may arise even if the TUR is not violated, since classical machines may not saturate the TUR in such relevant regimes. Hence in order to provide a fair and accurate assessment of practical quantum-thermodynamic advantages in thermal machines both optimization of the machine performance and comparison to classical models become necessary.
IV Classical-thermodynamic equivalents of thermal machines
@@
Throughout this paper, inspired by the notion of classical emulability introduced in Ref. [23], we use the term “classical-thermodynamic equivalent” (or simply “classical equivalent”) of a quantum thermal machine to refer to a thermal machine model with same bare Hamiltonian, $H_{0}$, but whose evolution can be described using only classical (Markovian) dynamics, while being capable of producing the same average currents as the quantum machine using the same amount of incoherent resources. Notice that if the classical equivalent machine is capable of producing the same average currents, this also ensures the same efficiency as the original quantum machine. However, quantum and classical equivalent models may different in general in their fluctuations.
@@
By using same incoherent resources we mean that the classical-thermodynamic equivalent model is in contact with the same thermal baths, and therefore have access to the same temperatures. Furthermore the requisite of having same bare Hamiltonian implies that the classical equivalent also has the same energy level structure as the original quantum machine, which allow us to identify them as the “same machine”. This is in contrast with other notions of classical analogs in thermal machines where all the parameters of the classical model are imposed to be exactly equal, including the couplings to the baths or the driving strength [9, 12] (this is also the case of introducing extra dephasing in the quantum model). Here instead we allow to vary these parameters as long as the requirements for weak-coupling and Markovian dynamics assumed for the thermal machine models are satisfied. This choice is not arbitrary, but allows for an stringer notion of quantum-thermodynamic advantage that avoids spurious “advantages” that may disappear by just slightly modifying some of the parameters in the model.
@@ IV.1 Hamiltonian-induced coherence
with transition rate
$\gamma_{\mathrm{uv}}^{cl}=\gamma_{\mathrm{vu}}^{cl}=\dfrac{4g^{2}}{\sum_{i}\left(\gamma_{\mathrm{u}i}+\gamma_{\mathrm{v}i}\right)}.$ (26)
@@ V Thermodynamic impact of coherence
We are now in a position to present our results that unveil the impact of steady-state coherence in the performance of quantum thermal machines by direct comparison with the corresponding classical equivalents as introduced above. We recall that, by construction, the classical equivalent machine reproduces the same average currents in all the transitions, and hence it has the same efficiency than the original machine. However, fluctuations in the output current (as captured by the variance), can differ in general between quantum and classical models, making the presence of coherence either beneficial or detrimental for the machine output reliability. In the following, we show that energetic coherence can lead to reliability improvements (for same output and efficiency), while, on the contrary, thermal machines operating with degenerate coherence can only perform equal or worst than their classical-equivalent counterparts.
@@
In order to address the impact of coherence in the reliability of the thermal machines, we introduce the fluctuations ratio $\mathcal{R}$ between quantum and classical machines as
$\mathcal{R}:=\dfrac{\mathrm{Var}[J_{\mathrm{out}}]}{\mathrm{Var}[J_{\mathrm{out}}^{\mathrm{cl}}]},$ (31)

### v2

# Reassessing quantum-thermodynamic enhancements in continuous thermal machines

@@ Abstract
Quantum coherence has been shown to impact the operational capabilities of quantum systems performing thermodynamic tasks in a significant way, and yet the possibility and conditions for genuine coherence-enhanced thermodynamic operation remain unclear. We show that for steady-state quantum thermal machines —both autonomous and externally driven— that interact weakly with thermal reservoirs and work sources, the presence of coherence induced by perturbations in the machine Hamiltonian guarantees a genuine thermodynamic advantage. Such advantage applies both for the cases in which the induced coherence is between levels with different energies or between degenerate levels. On the other hand, we show that engines subjected to noise-induced coherence, can be outperformed by classical stochastic engines using exactly the same set of (incoherent) resources. We illustrate our results with three prototypical models of heat engines and refrigerators and employ multi-objective optimization techniques to characterize quantum-enhanced regimes in connection with the thermodynamic uncertainty relation and beyond it.
@@ I Introduction
Figure 1: Schematic representation of three quantum thermal machine models. (a) The coherent three-level amplifier with couplings to baths at hot $\beta_{\mathrm{h}}$(red) and cold $\beta_{\mathrm{c}}$(blue) temperatures, as well as coherent external driving (green thunderbolt), (b) the three-qubit autonomous (absorption) refrigerator where each qubit is locally coupled to baths at hot (red), cold (blue), and intermediate (yellow) temperatures $\beta_{\mathrm{m}}$, and (c) the noise-induced-coherence machine showing collective jumps induced by the baths at hot (red) and cold (blue) temperatures, together with a classical work source given by an infinite-temperature bath (green). Plain double arrows represent jumps between the machine energy levels induced by the baths while wavy arrows represent coherent interactions. Collective jumps in (c) are represented by triple arrows.
@@
In this paper we show how to identify genuine coherence-induced quantum thermodynamic advantage and how to quantitatively characterize it in steady-state quantum thermal machines by combining three main ingredients. The first one is the systematic construction of thermodynamically-equivalent classical thermal machines that are able to produce the same average currents than their quantum-coherent counterparts, while using exactly the same amount of incoherent resources (essentially bath temperatures and energy structure). However, even if same average currents can be reproduced by classical models, fluctuations as captured by the variance of the currents might present significant differences. The comparison of fluctuations in the output currents (reliability of the machine) with respect to the classical counterpart is henceforth the second necessary ingredient in our analysis. This is connected with violations in the so-called thermodynamic uncertainty relation (TUR) [24, 25, 26], which provides a universal trade-off relation between power, efficiency, and output reliability for any classical Markovian thermal machine operating at steady-state conditions [27]. The TUR sets up a model-independent limit in the maximum reliability achievable by any classical machine at a given power output and efficiency. Hence the observation of TUR violations in quantum Markovian machines working in steady-state conditions [28, 29, 30, 31, 32, 33] may be considered as an unambigüous witness of a quantum thermodynamic signature. However the contrary is not necessarily true, i.e. the absence of a TUR violation does not imply the absence of thermodynamic enhancements in the machine, as we explicitly show for different classes of quantum machines, like e.g. autonomous multipartite machines. Moreover, the violation of the TUR in specific parameter regimes is not yet a sufficient (nor necessary) condition for ensuring a practical quantum-thermodynamic advantage in thermal machines, since these parameter regimes may be far from optimal performance and therefore be useless after all. The third ingredient that we employ here is then the use of multi-objective optimization techniques (Pareto optimization) [34, 35, 36, 37, 38, 39] to actually confirm enhanced stability (precision), power output or efficiency, with respect to classical thermodynamically-equivalent machines, in relevant (optimal) regimes of operation of the machine, such as maximum power conditions.
@@
Our general results show systematic quantum-thermodynamic advantages in the presence of Hamiltonian-induced coherence that are applicable to a broad class of steady-state quantum thermal machines (both autonomous or externally driven) in the weak coupling limit. Moreover, we illustrate our results by employing three well-known (prototypical) models showing coherent-induced evolution (see Fig. 1), each corresponding to one of the three possible types of coherence that can arise in the working medium when approaching the steady state: (a) coherence between different energy levels (energetic coherence) induced by external driving, (b) coherence between levels with degenerate energies induced by internal Hamiltonian couplings, and (c) noise-induced coherence on degenerate levels induced by the reservoir. In the case of Hamiltonian-induced coherence (cases a and b) practical quantum-thermodynamic advantages are unambiguously identified, leading to enhanced optimal regimes not allowed by any equivalent classical engine using the same amount of thermodynamic resources. On the contrary, noise-induced coherence between degenerate levels (case c) can lead to disadvantages in the performance, even in cases where the quantum machine dynamics contains intrinsic quantum features.
@@
This paper is organized as follows: Sec. II introduces the different thermal machine models in detail. Sec. III discusses how to characterize the thermodynamic performance in such systems. In Sec. IV, we describe the concept of classical equivalent machines and provide a general recipe for constructing them in the presence of different types of coherence. In Sec. V, we show that in the presence of Hamiltonian coherence, quantum improvements over classical equivalent machines can always be achieved for a broad class of quantum models including unicyclic engines and multipartite (multicycle) engines, whereas in the presence of noise-induced coherence, the performance of quantum machines over their classical counterparts depends on the system parameters. In Sec. VI we verify our general statements using three specific models of quantum thermal machines that use different types of coherence to operate. In Sec. VII we optimize both quantum and classical thermal machines and compare their optimal solutions. Finally, Sec. VIII provides a summary and conclusions of our main results.
@@ II Quantum thermal machines models
where $H(t)$ is given in Eq. (1) and $\mathcal{D}^{(r)}_{k}(\rho)$ denote the so-called dissipators, taking into account the effects of dissipative process $k$ from thermal reservoir $r$ on the system (we take $\hbar=1$ through the paper). These dissipators are given in terms of Lindblad operators $L^{(r)}_{k}$ associated to each reservoir:
@@
with $\alpha_{ij}^{k}=1$ if $\epsilon_{j}-\epsilon_{i}=\Delta\epsilon_{k}$ and 00 otherwise. In the case of non-degenerate transitions, the above operators reduce to simple jumps $L_{k}^{(r)}=\sqrt{\gamma_{ij}}\ket{j}\bra{i}$ | between energy levels $i\rightarrow j$. However, the operators in Eq. (4) can also describe collective jumps where two or more transitions with same (degenerate) gap $\Delta\epsilon_{k}$ in $H_{0}$ may occur simultaneously, e.g. $i\rightarrow j$ and $i\rightarrow j^{\prime}$ if $\epsilon_{j}=\epsilon_{j}^{\prime}$. In any case, every transition is connected to a single thermal bath, the rates $\gamma_{ij}\geq 0$ being time-independent and verifying the local detailed balance relation $\gamma_{ij}=\gamma_{ji}~{}e^{-\beta_{r}(\epsilon_{j}-\epsilon_{i})}$, with $\beta_{r}=1/{k_{B}T_{r}}$ the inverse temperature of the bath $r$ and $k_{B}$ Boltzmann’s constant. In the long-time limit, the evolution dictated by Eq. (3) converges to a steady-state, verifying $\mathcal{L}(\pi)=0$, where we denote $\pi(t)$ the steady state density operator. We notice that, due to the presence of the time-dependent Hamiltonian $V(t)$, the steady state can show a (residual) periodic time-dependence in the phase in the Schrödinger picture.
Although the results we present apply to any model of quantum thermal machine whose evolution can be described within the general framework introduced above, we will often particularize to three representative and well-known models of quantum thermal machines, as illustrated in Fig. 1. They lead to three different types of coherent evolution respectively: the coherent three-level maser (Fig. 1a) which induces energetic coherence in the $H_{0}$ basis; the three-qubit autonomous quantum refrigerator (Fig. 1b), leading to Hamiltonian-induced coherence in a degenerate subspace of $H_{0}$, and the noise-induced-coherence engine (Fig. 1c), showing coherence between degenerate levels induced by collective bath transitions.
@@ II.3 Noise-induced-coherence machine
A notable example of this coherence is found in a 4-level absorption refrigerator than can also work as a heat engine, whose thermodynamics have been examined in previous studies [66, 69, 68]. In this system, two of the levels possess the same energy, as depicted in Fig. 1c. These levels are subjected to the influence of two distinct thermal baths, whose action results in the emergence of degenerate coherence that persist in the system’s steady state. The system’s Hamiltonian in this case is as follows
@@ III Thermodynamic performance
where we recall that $\pi(t)$ may acquire a periodic time-dependence (in Schrödinger picture) due to the presence of non-diagonal elements (coherences) in the steady-state density operator. We also emphasize that for weak perturbations as the ones considered here, only the bare Hamiltonian $H_{0}$ enters in the heat currents [71, 72], ensuring consistency with the second law [73] (see also Appendix A). The first law takes the form $\langle\dot{W}\rangle=\sum_{r}\langle\dot{Q}_{r}\rangle$, imposing that any output power is sustained by input heat currents from the baths. Explicit expressions of the heat currents $\langle\dot{Q}_{r}\rangle$ valid for generic machines (with or without degeneracy) are given in Appendix A.
@@
where $\langle\dot{N}_{r}\rangle$ is the probability current associated to the reservoir $r$ (flux of quanta). It results from the fact that in the models, baths are associated to a single energy gap $\Delta\epsilon_{r}$. In the steady state, due to the absence of leaks, the system will exchange excitations with all baths at the same rate, i.e. $\forall i,j$ we have $\langle\dot{N}_{i}\rangle=\langle\dot{N}_{j}\rangle:=\langle\dot{N}\rangle$. In addition, the first law constrains the average value of the work to satisfy $\langle\dot{W}\rangle=\sum_{r}\Delta\epsilon_{r}\langle\dot{N}\rangle=\omega_{d}\langle\dot{N}\rangle$. The ratio between the different steady-state currents hence verify
@@
By combining the above proportionality relations with the efficiency bounds above, we can construct a table for the operational modes of each model (see Table 1).
@@
In any of the three cases, the TUR implies that beyond a certain threshold, a classical Markovian engine can only enhance its precision in the output (cooling) power at the cost of either reducing the output itself or reducing the energy efficiency, so that the above ratio remains bounded by $2$, i.e. $\mathcal{Q}\geq 2$. However, some models of quantum thermal machines have been shown to produce violations of the TUR, that is, they verify $\mathcal{Q}<2$ (see e.g. Refs. [28, 29, 30, 31, 32, 33]). Such violations act as a witness indicating an enhanced tradeoff between power, precision and efficiency that arise in certain parameter regimes. Nevertheless, it is in principle unclear whether that violations may occur in relevant regimes, i.e. where the machine performs in an optimal way, providing a practical quantum-thermodynamic advantage. In addition, such quantum-thermodynamic advantages may arise even if the TUR is not violated, since classical machines may not saturate the TUR in such relevant regimes. Hence in order to provide a fair and accurate assessment of quantum-thermodynamic advantages in thermal machines both optimization of the machine performance and comparison to classical models become necessary.
IV Classical thermodynamic equivalents of thermal machines
@@
Throughout this paper, inspired by the notion of classical emulability introduced in Ref. [23], we use the term “classical thermodynamic-equivalent” (or simply “classical equivalent”) of a quantum thermal machine to refer to a thermal machine model with same bare Hamiltonian, $H_{0}$, but whose evolution can be described using only classical Markovian dynamics (stochastic jumps between the energy levels), while being capable of producing the same average currents as the quantum machine, by using the same amount of incoherent resources. Notice that if the classical equivalent machine is capable of producing the same average currents, this ensures the same power output and efficiency as the original quantum machine. However, quantum and classical equivalent models may different in general in their fluctuations.
@@
By using same incoherent resources we mean that the classical-thermodynamic equivalent model is in contact with the same thermal baths, and therefore have access to the same temperatures. Furthermore, the requisite of having same bare Hamiltonian $H_{0}$ implies that the classical equivalent also has the same energy level structure as the original quantum machine, which allow us to identify them as the “same machine”. This is in contrast with other notions of classical analogs in thermal machines where all the parameters of the classical model are imposed to be exactly equal, including the couplings to the baths or the driving strength [9, 12] (this is also the case of introducing extra dephasing in the quantum model). Here instead we allow to vary these parameters as long as the requirements for weak-coupling and Markovian dynamics assumed for the thermal machine models are satisfied. This choice is not arbitrary, but allows for an stringer notion of genuine quantum-thermodynamic advantage that avoids spurious “advantages” that may disappear by just slightly modifying some of the parameters in the classical model.
@@ IV.1 Hamiltonian-induced coherence
with transition rates
$\gamma_{\mathrm{uv}}^{\rm cl}=\gamma_{\mathrm{vu}}^{\rm cl}=\dfrac{4g^{2}}{\sum_{i}\left(\gamma_{\mathrm{u}i}+\gamma_{\mathrm{v}i}\right)}.$ (26)
@@ V Thermodynamic impact of coherence
We are now in a position to present our general results that unveil the impact of steady-state coherence in the performance of quantum thermal machines by direct comparison with the classical thermodynamic-equivalent machine introduced above. We recall that, by construction, the classical equivalent machine reproduces the same average currents in all the transitions, and hence it has the same efficiency than the original machine. However, fluctuations in the output current (as captured by the variance), can differ in general between quantum and classical models, making the presence of coherence either beneficial or detrimental for the machine output reliability. In the following, we show that Hamiltonian coherence always lead to reliability improvements (for same output and efficiency), while thermal machines operating with noise-induced coherence can either perform better or worse than their classical-equivalent counterparts.
@@
In order to address the impact of coherence in the reliability of the thermal machines, we introduce the relative difference in the fluctuations between classical and quantum machines as
$\mathcal{R}:=\dfrac{\mathrm{Var}[J_{\mathrm{out}}^{\mathrm{cl}}]-\mathrm{Var}[J_{\mathrm{out}}]}{\mathrm{Var}[J_{\mathrm{out}}]},$ (31)